# Data-Driven Nonconvex Reachability Analysis using Exact Multiplication

### v1

@@ III-C Iterative Refinement of Set of Models
Similar to Lemma 1, given any input-state data $D_{i}=(X_{i}^{+},X_{i}^{-},U_{i}^{-})$ of system (1) over the interval $t\in[T_{i-1},T_{i}]$, for $i\geq 1$, and that Assumption 2 holds, the resulting MZ $\mathcal{M}_{i}^{\Sigma}$ contains all matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ consistent with the data and noise bound, i.e., $\mathcal{N}_{\Sigma}\subseteq\mathcal{M}_{i}^{\Sigma}$. At iteration $i$ of the refinement phase ($i\geq 1$), we intersect the set of models ${\mathcal{M}}_{i}^{\Sigma}$ derived from newly obtained input-state trajectories with the previous model set ^$\hat{\mathcal{M}}_{i-1}^{\Sigma}$ to get refined set ^$\hat{\mathcal{M}}_{i}^{\Sigma}$. The initial set of models and refined sets are subsequently used to compute an over-approximation of the reachable sets of the unknown system, as illustrated in Fig. 1.
@@
^$\hat{\mathcal{M}}_{i}^{\Sigma}=\mathcal{M}_{i}^{\Sigma}\cap\hat{\mathcal{M}}_{i-1}^{\Sigma},\quad i\geq 1,\quad\text{with}\quad\hat{\mathcal{M}}_{0}^{\Sigma}=\mathcal{M}_{0}^{\Sigma}$ (15)
@@ III-D Exact Multiplication for Non-Convex Reachability
At time $k$, as depicted in Fig. 1, we need to iteratively multiply ^$\hat{\mathcal{M}}_{i}^{\Sigma}$ with the cartesian product of reachable set ^$\hat{\mathcal{R}}_{k}$ and the input set $\mathcal{U}_{k}$ to obtain the next reachable set ^$\hat{\mathcal{R}}_{k+1}$. Although ^$\hat{\mathcal{M}}_{i}^{\Sigma}$ is a CMZ, notice that CMZs are a special case CPMZ. Therefore, to be more general, we provide an exact multiplication method for CPMZ with CPZ.
@@
${\texttt{mergeID}}(\mathcal{Y},\mathcal{P})\!=\!\big{\{}\!\underbrace{\langle C_{\mathcal{Y}},\!G_{\mathcal{Y}},\!\overline{E}_{\mathcal{Y}},\!A_{\mathcal{Y}},\!B_{\mathcal{Y}},\!\overline{R}_{\mathcal{Y}},\!\mathsf{id}_{\mathcal{YP}}\rangle_{\text{CPMZ}}}_{\bar{\mathcal{Y}}},\\ \underbrace{\langle c_{\mathcal{P}},\!G_{\mathcal{P}},\!\overline{E}_{\mathcal{P}},\!A_{\mathcal{P}},\!b_{\mathcal{P}},\!\overline{R}_{\mathcal{P}},\!\mathsf{id}_{\mathcal{YP}}\rangle_{\text{CPZ}}}_{\bar{\mathcal{P}}}\big{\}}.$ (16)
@@
$g_{f}^{(k)}=G^{(i)}_{\mathcal{Y}}G_{\mathcal{P}}^{(\cdot,j)},\quad k=h_{\mathcal{P}}(i-1)+j$
for $i=1,\dots,\gamma$ and $j=1,\dots,h_{\mathcal{P}}$.
@@
$\exists\hat{\alpha}_{([1:a])}$ ::: $Y$ $=C_{\mathcal{Y}}+\sum_{i=1}^{\gamma}\bigg{(}\prod_{k=1}^{a}\hat{\alpha}_{(k)}^{\overline{E}_{\mathcal{Y}}^{(k,i)}}\bigg{)}G_{\mathcal{Y}}^{(i)}$
$\exists\hat{\alpha}_{([1:a])}$ ::: $p$ $=\!c_{\mathcal{P}}+\sum_{i=1}^{h_{\mathcal{P}}}\bigg{(}\prod_{k=1}^{a}\hat{\alpha}_{(k)}^{\overline{E}_{\mathcal{P}}^{(k,i)}}\bigg{)}G_{\mathcal{P}}^{(\cdot,i)}$ (18)
where $a=|\mathcal{H}|+\gamma$, =$\mathcal{H}=\left\{i\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathsf{id}_{{\mathcal{P}}(i)}\not\in\mathsf{id}_{\mathcal{Y}}\right\}$. Thus, we have
@@
Output: Reachable sets ^$\hat{\mathcal{R}}_{k}$ for $k\in\mathbb{N}$.
@@
1:^$\hat{\mathcal{R}}_{0}\leftarrow\mathcal{X}_{0}$
2:$\mathcal{M}_{0}^{\Sigma}\leftarrow(X_{0}^{+}-\mathcal{M}_{w})\begin{bmatrix}X_{0}^{-}\\ U_{0}^{-}\end{bmatrix}^{\dagger}$
3:Set ^$\hat{\mathcal{M}}^{\Sigma}\leftarrow\mathcal{M}_{0}^{\Sigma}$
4:Set $X^{+}\leftarrow[\leavevmode\nobreak\ ]$, $X^{-}\leftarrow[\leavevmode\nobreak\ ]$, $U^{-}\leftarrow[\leavevmode\nobreak\ ]$
@@
2:     Construct online data matrices $X^{+}=[X^{+}\leavevmode\nobreak\ x_{(k)}]$, $X^{-}=[X^{-}\leavevmode\nobreak\ x_{(k-1)}]$, $U^{-}=[U^{-}\leavevmode\nobreak\ u_{(k-1)}]$
@@
5:         ^$\hat{\mathcal{M}}^{\Sigma}\leftarrow\mathcal{M}^{\Sigma}\cap\hat{\mathcal{M}}^{\Sigma}$
6:         $X^{+}\leftarrow[\leavevmode\nobreak\ ]$, $X^{-}\leftarrow[\leavevmode\nobreak\ ]$, $U^{-}\leftarrow[\leavevmode\nobreak\ ]$
@@
8:     ^$\hat{\mathcal{R}}_{k+1}\leftarrow\hat{\mathcal{M}}^{\Sigma}\otimes(\hat{\mathcal{R}}_{k}\times\mathcal{U}_{k})\boxplus\mathcal{Z}_{w}$
@@
Figure 4: Comparative Projection of Reachable Sets from Input-State Data Using Algorithm 1 and Alanwar et al. [3]
@@ IV Numerical simulations
Consider a five-dimensional system which is a discretization of the system used in [3] with sampling time $0.05$ sec. All used code to reproduce our results are publicly available [^1] https://github.com/TUM-CPS-HN/Data-Driven-Nonconvex-Reachability-Analysis. The system has the following matrices.
@@
In the first experiment, the initial non-convex set is chosen to be $\mathcal{X}_{0}=\langle c_{0},G_{0},E_{0},[\leavevmode\nobreak\ ],[\leavevmode\nobreak\ ],[\leavevmode\nobreak\ ],\mathsf{id}_{0}\rangle_{\mathcal{P}}$ , where
$\begin{split}c_{0}=1_{5\times 1},\leavevmode\nobreak\ G_{0}=0.1I_{5},\leavevmode\nobreak\ E_{0}&=\begin{bmatrix}2&1&0&0&0\\ 1&2&0&0&0\\ 0&0&2&1&0\\ 0&0&1&2&1\\ 0&0&0&1&2\end{bmatrix}.\end{split}$
@@
For the second experiment, a convex initial set $\hat{\mathcal{X}}_{0}=\langle 1_{5\times 1},0.1I_{5}\rangle$ is used to compare reachable sets computed via Algorithm 1 and the method of Alanwar et al. [3]. In both experiments, the input set is defined as $\mathcal{U}_{k}=\langle 10,0.25\rangle$, and random noise is sampled from the zonotope $\mathcal{Z}_{w}=\langle 0,[0.005,\dots,0.005]^{\top}\rangle$.
@@
Firstly, the reachable sets obtained from a non-convex initial set using both the nominal system model and Algorithm 1 are illustrated in Fig. 3. The reachable set ^$\hat{\mathcal{R}}_{1}$ corresponds to the one computed using the model set $\mathcal{M}_{1}^{\Sigma}$, without applying any set refinement. $\hat{\mathcal{R}}$ represents the reachable set computed after performing the first online set refinement using new incoming data through Algorithm 1. Owing to the exact set multiplication employed in the algorithm, the non-convexity of the initial set is preserved throughout the reachability propagation process. Furthermore, Proposition 4 guarantees that the initial reachability computation remains applicable to non-convex initial sets by employing exact multiplication, thus addressing the limitation of the method proposed by Alanwar et al. [3], which is confined to convex cases.
@@
Secondly, to evaluate the performance of the proposed algorithm, a comparative analysis was conducted under two scenarios. First, trajectory data collected over the interval $[0,T_{1}]$ with $N_{1}$ trajectories were used to perform the offline initialization step of Algorithm 1. At time $T_{1}$, new data over $[T_{1},T_{2}]$ containing $N_{2}$ trajectories became available, triggering the online set refinement step of Algorithm 1 and yielding the updated reachable set $\hat{\mathcal{R}}$. For consistency and fair comparison with the method of Alanwar et al. [3], we set $N_{1}=N_{2}$. The reachable set ^$\hat{\mathcal{R}}_{a}$ computed using the method of Alanwar et al. [3] is based on the combined dataset $D_{a}$ over the full interval $[0,T_{2}]$.
@@
As illustrated in Fig. 4, the reachable set $\hat{\mathcal{R}}$ computed using the Algorithm 1 is less conservative than ^$\hat{\mathcal{R}}_{a}$ obtained using the method of Alanwar et al. [3]. This improvement stems from the incremental refinement of the model set, where newly acquired data are used to further constrain the previously computed set of models. As a result, a more accurate representation of system behavior is achieved in the form of a CMZ, rather than a MZ. The use of exact set multiplication between CMZ and CPZ in Algorithm 1 ensures exact set propagation, thereby further reducing conservativeness. These results highlight the advantages of the proposed incremental approach over direct one-shot data-driven methods, demonstrating its improved ability to capture system dynamics and perform set propagation.
@@ V Conclusion
We presented a data-driven framework for reachability analysis using non-convex set representations. We introduced a set refinement strategy that incrementally incorporates newly collected input-state data, enabling iterative model set updates. This refinement process is supported by an intersection operation for CMZs, allowing system dynamics to be represented more accurately over time. Furthermore, we introduced the CPMZ as a new set representation that enables exact multiplication with CPZs. Since CMZs are a special case of CPMZs, the proposed method seamlessly applies to the exact multiplication between CMZs and CPZs. This enables the exact propagation of non-convex reachable sets without introducing over-approximation. The proposed approach thus achieves improved accuracy and reduced conservativeness in safety verification. Future work will focus on extending the framework to systems with nonlinear dynamics.

### v2

@@ III-C Iterative Refinement of Set of Models
Similar to Lemma 1, given any input-state data $D_{i}=(X_{i}^{+},X_{i}^{-},U_{i}^{-})$ of system (1) over the interval $t\in[T_{i-1},T_{i}]$, for $i\geq 1$, and that Assumption 2 holds, the resulting MZ $\mathcal{M}_{i}^{\Sigma}$ contains all matrices $\begin{bmatrix}\Phi&\Gamma\end{bmatrix}$ consistent with the data and noise bound, i.e., $\mathcal{N}_{\Sigma}\subseteq\mathcal{M}_{i}^{\Sigma}$. At iteration $i$ of the refinement phase ($i\geq 1$), we intersect the set of models ${\mathcal{M}}_{i}^{\Sigma}$ derived from newly obtained input-state trajectories with the previous model set $\hat{\mathcal{M}}_{i-1}^{\Sigma}$ to get refined set $\hat{\mathcal{M}}_{i}^{\Sigma}$. The initial set of models and refined sets are subsequently used to compute an over-approximation of the reachable sets of the unknown system, as illustrated in Fig. 1.
@@
$\hat{\mathcal{M}}_{i}^{\Sigma}=\mathcal{M}_{i}^{\Sigma}\cap\hat{\mathcal{M}}_{i-1}^{\Sigma},\quad i\geq 1,\quad\text{with}\quad\hat{\mathcal{M}}_{0}^{\Sigma}=\mathcal{M}_{0}^{\Sigma}$ (15)
@@ III-D Exact Multiplication for Non-Convex Reachability
At time $k$, as depicted in Fig. 1, we need to iteratively multiply $\hat{\mathcal{M}}_{i}^{\Sigma}$ with the cartesian product of reachable set $\hat{\mathcal{R}}_{k}$ and the input set $\mathcal{U}_{k}$ to obtain the next reachable set $\hat{\mathcal{R}}_{k+1}$. Although $\hat{\mathcal{M}}_{i}^{\Sigma}$ is a CMZ, notice that CMZs are a special case CPMZ. Therefore, to be more general, we provide an exact multiplication method for CPMZ with CPZ.
@@
$\hskip-7.0pt{\texttt{mergeID}}(\mathcal{Y},\mathcal{P})\!=\!\big\{\!\underbrace{\langle C_{\mathcal{Y}},\!G_{\mathcal{Y}},\!\overline{E}_{\mathcal{Y}},\!A_{\mathcal{Y}},\!B_{\mathcal{Y}},\!\overline{R}_{\mathcal{Y}},\!\mathsf{id}_{\mathcal{YP}}\rangle_{\text{CPMZ}}}_{\bar{\mathcal{Y}}},\\ \underbrace{\langle c_{\mathcal{P}},\!G_{\mathcal{P}},\!\overline{E}_{\mathcal{P}},\!A_{\mathcal{P}},\!b_{\mathcal{P}},\!\overline{R}_{\mathcal{P}},\!\mathsf{id}_{\mathcal{YP}}\rangle_{\text{CPZ}}}_{\bar{\mathcal{P}}}\big\}.$ (16)
@@
$g_{f}^{(k)}=G^{(i)}_{\mathcal{Y}}G_{\mathcal{P}}^{(\cdot,j)},\quad k=h_{\mathcal{P}}(i-1)+j,$
for $i=1,\dots,\gamma$, $j=1,\dots,h_{\mathcal{P}}$.
@@
$\exists\hat{\alpha}_{([1:a])}$ $:$ $Y$ $=C_{\mathcal{Y}}+\sum_{i=1}^{\gamma}\bigg(\prod_{k=1}^{a}\hat{\alpha}_{(k)}^{\overline{E}_{\mathcal{Y}}^{(k,i)}}\bigg)G_{\mathcal{Y}}^{(i)}$
$\exists\hat{\alpha}_{([1:a])}$ $:$ $p$ $=\!c_{\mathcal{P}}+\sum_{i=1}^{h_{\mathcal{P}}}\bigg(\prod_{k=1}^{a}\hat{\alpha}_{(k)}^{\overline{E}_{\mathcal{P}}^{(k,i)}}\bigg)G_{\mathcal{P}}^{(\cdot,i)}$ (18)
where $a=|\mathcal{H}|+\gamma$, $\mathcal{H}=\left\{i\penalty 10000\ |\penalty 10000\ \mathsf{id}_{{\mathcal{P}}(i)}\not\in\mathsf{id}_{\mathcal{Y}}\right\}$. Thus, we have
@@
Output: Reachable sets $\hat{\mathcal{R}}_{k}$ for $k\in\mathbb{N}$.
@@
[…]
3:Set $\hat{\mathcal{M}}^{\Sigma}\leftarrow\mathcal{M}_{0}^{\Sigma}$
[…]
@@
2:  Construct online data matrices $^{+}=[X^{+}\penalty 10000\ x_{(k)}], X=[X$$^{-}=[X^{-}\penalty 10000\ x_{(k-1)}$], $U^{-}=[U^{-}\penalty 10000\ u_{(k-1)}]$
@@
5:   $\hat{\mathcal{M}}^{\Sigma}\leftarrow\mathcal{M}^{\Sigma}\cap\hat{\mathcal{M}}^{\Sigma}$
[…]
@@
8:  $\hat{\mathcal{R}}_{k+1}\leftarrow\hat{\mathcal{M}}^{\Sigma}\otimes(\hat{\mathcal{R}}_{k}\times\mathcal{U}_{k})\boxplus\mathcal{Z}_{w}$
@@
Figure 4: Comparative Projection of Reachable Sets from Input-State Data Using Algorithm 1 and Alanwar et al. [Alanwar2023Datadriven]
@@ IV Numerical simulations
Consider a five-dimensional system which is a discretization of the system used in [Alanwar2023Datadriven] with sampling time $0.05$ sec. All used code to reproduce our results are publicly available [^1] https://github.com/TUM-CPS-HN/Data-Driven-Nonconvex-Reachability-Analysis. The system has the following matrices.
@@
In the first experiment, the initial non-convex set is chosen to be $\mathcal{X}_{0}=\langle c_{0},G_{0},E_{0},[\penalty 10000\ ],[\penalty 10000\ ],[\penalty 10000\ ],\mathsf{id}_{0}\rangle_{\mathcal{P}}$ , where
[…]
@@
For the second experiment, a convex initial set $\hat{\mathcal{X}}_{0}=\langle 1_{5\times 1},0.1I_{5}\rangle$ is used to compare reachable sets computed via Algorithm 1 and the method of Alanwar et al. [Alanwar2023Datadriven]. In both experiments, the input set is defined as $\mathcal{U}_{k}=\langle 10,0.25\rangle$, and random noise is sampled from the zonotope $\mathcal{Z}_{w}=\langle 0,[0.005,\dots,0.005]^{\top}\rangle$.
@@
Firstly, the reachable sets obtained from a non-convex initial set using both the nominal system model and Algorithm 1 are illustrated in Fig. 3. The reachable set $\tilde{\mathcal{R}}_{k}$ corresponds to the one computed using the model set $\mathcal{M}_{0}^{\Sigma}$, without applying any set refinement. $\hat{\mathcal{R}}_{k}$ represents the reachable set computed after performing the first online set refinement using new incoming data through Algorithm 1. Owing to the exact set multiplication employed in the algorithm, the non-convexity of the initial set is preserved throughout the reachability propagation process. Furthermore, Proposition 4 guarantees that the initial reachability computation remains applicable to non-convex initial sets by employing exact multiplication, thus addressing the limitation of the method proposed by Alanwar et al. [Alanwar2023Datadriven], which is confined to convex cases.
@@
Secondly, to evaluate the performance of the proposed algorithm, a comparative analysis was conducted under two scenarios. First, trajectory data collected over the interval $0,T_{1}]$ with $N_{1}$ trajectories were used to perform the offline initialization step of Algorithm 1. At time $T_{1}$, new data over $T_{1},T_{2}]$ containing $N_{2}$ trajectories became available, triggering the online set refinement step of Algorithm 1 and yielding the updated reachable set $\hat{\mathcal{R}}_{k}$. For consistency and fair comparison with the method of Alanwar et al. [Alanwar2023Datadriven], we set $N_{1}=N_{2}$. The reachable set $\bar{\mathcal{R}}_{k}$ computed using the method of Alanwar et al. [Alanwar2023Datadriven] is based on the combined dataset $D_{a}$ over the full interval $0,T_{2}]$.
@@
As illustrated in Fig. 4, the reachable set $\hat{\mathcal{R}}_{k}$ computed using the Algorithm 1 is less conservative than $\bar{\mathcal{R}}_{k}$ obtained using the method of Alanwar et al. [Alanwar2023Datadriven]. This improvement stems from the incremental refinement of the model set, where newly acquired data are used to further constrain the previously computed set of models. As a result, a more accurate representation of system behavior is achieved in the form of a CMZ, rather than a MZ. The use of exact set multiplication between CMZ and CPZ in Algorithm 1 ensures exact set propagation, thereby further reducing conservativeness. These results highlight the advantages of the proposed incremental approach over direct one-shot data-driven methods, demonstrating its improved ability to capture system dynamics and perform set propagation.
@@ V Conclusion
We presented a data-driven framework for reachability analysis using non-convex set representations. We introduced a set refinement strategy that incrementally incorporates newly collected input-state data, enabling iterative model set updates. This refinement process is supported by an intersection operation for CMZs, allowing system dynamics to be represented more accurately over time. Furthermore, we introduced the CPMZ as a new set representation that enables exact multiplication with CPZs. Since CMZs are a special case of CPMZs, the proposed method seamlessly applies to the exact multiplication between CMZs and CPZs. This enables the exact propagation of non-convex reachable sets by avoiding the over-approximation errors inherent in the non-exact arithmetic of zonotope operations. The proposed approach thus achieves improved accuracy and reduced conservativeness in safety verification. Future work will focus on extending the framework to systems with nonlinear dynamics.